OpenStudy (anonymous):

if y varies directly as x and y=12 when x=8, find y when x= 14. Please help me with this. i have no idea how to do it

OpenStudy (diyadiya):

y varies directly with x y=kx 12=k*8 therefore k=12/8=3/2 NOw when x=14 (substitute the value of k) y=kx y=(3/2)*14 y= .. ?

OpenStudy (diyadiya):

Did you understand ?

OpenStudy (anonymous):

so u divide 3 and 2 and then times that by 14? i have like 8 problems that deal with this and ive done them all wrong....do u mind helping me with them?

OpenStudy (diyadiya):

Sure ,but you've to remember this When it varies directly we have to use y=kx or x=ky (according to the question ) when it varies inversely x=k*(1/y) or y=k(1/x) (according to the question)

OpenStudy (anonymous):

so then what the anwser would be is 14?

OpenStudy (diyadiya):

No we got y= 3/2 * 14

OpenStudy (anonymous):

so 21

OpenStudy (diyadiya):

Right!

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

Will you help me with the others? please and thank you

OpenStudy (diyadiya):

Sure :) Post it

OpenStudy (anonymous):

suppose y varies jointly as x and z. Find y when x=9 and z=-3, if y=-50 when z is 5 and x is -10

OpenStudy (diyadiya):

Ok so according to the question y=(k)xz Now you need to find the constant k

OpenStudy (diyadiya):

@RiyaMoon put the values of y , x & z and find k

OpenStudy (anonymous):

-44? i really understand this at all....

OpenStudy (diyadiya):

In the question its given y varies jointly as x and z that means y is directy proportional so we take the equation as y=(k)xz do you understand this ?

OpenStudy (anonymous):

Not really... weres K?

OpenStudy (diyadiya):

K is a constant of variation.

OpenStudy (diyadiya):

YOu have to remember this when y varies jointly/directly with x y=kx when x varies jointly/directly with y x=ky got it ?

OpenStudy (anonymous):

some what yea

OpenStudy (diyadiya):

" y=-50 when z is 5 and x is -10" here y ,x & z are given so you can easily find out k using the equation y=(k)xz and use the value of k to find out y when x=9 and z=-3

OpenStudy (anonymous):

-50=(k) 9 x -3?

OpenStudy (diyadiya):

Nope read the question again "Find y when x=9 and z=-3" So you have to use this x & z to find out y in the end now we have to find out k using y=-50 z= 5 and x = -10

OpenStudy (anonymous):

so then 5/-10? this is confusing me even more...

OpenStudy (diyadiya):

y=(k)xz got this ?

OpenStudy (anonymous):

yes but it doesnt help that theres so many numbers...

OpenStudy (diyadiya):

Its very easy dont worry :)

OpenStudy (diyadiya):

you should be careful while choosing the numbers

OpenStudy (anonymous):

Lol it might be easy for someone who likes math, math has always been my down fall

OpenStudy (diyadiya):

"Find y when x=9 and z=-3' Now you know the equation y=(k)xz so can you find y with this ? No because you have two unknown variables y & k

OpenStudy (diyadiya):

So we move on to the next part of the question , where all the values are given " if y=-50 when z is 5 and x is -10" Now can you use this to find out k ? (Yes ,using the equation y= (k) xz

OpenStudy (anonymous):

How do i plug them in? i just did this and u said i did it wrong

OpenStudy (diyadiya):

y=(k)xz -50=(k)5*(-10)

OpenStudy (anonymous):

so you do 5 * (-10)? first ?

OpenStudy (diyadiya):

Yes

OpenStudy (anonymous):

-50

OpenStudy (diyadiya):

Right :) Now what's k ?

OpenStudy (anonymous):

-50?

OpenStudy (diyadiya):

Nope you're doing it wrong y=(k)xz -50= (k)*5(-10) -50=k*(-50) Now what's k?

OpenStudy (anonymous):

-50k?

OpenStudy (diyadiya):

Nope -50=k(-50) Divide both sides by -50 now what do you get?

OpenStudy (anonymous):

k=1

OpenStudy (diyadiya):

Right :D

OpenStudy (diyadiya):

Now can you find y when x=9 and z=-3

OpenStudy (anonymous):

so 1= 9/-3?

OpenStudy (diyadiya):

where is y ?

OpenStudy (anonymous):

y isnt found yet. so would it be 9/-3*1?

OpenStudy (diyadiya):

Nope why do you have to divide 9/-3 ? remember the equation y=(k)xz

OpenStudy (anonymous):

urg this is so confusing.....

OpenStudy (anonymous):

so then itss 1*9*-3?

OpenStudy (diyadiya):

Yes Exactly :)

OpenStudy (diyadiya):

So what is y?

OpenStudy (anonymous):

-27

OpenStudy (diyadiya):

Right ! :)

OpenStudy (anonymous):

kk thanks

OpenStudy (diyadiya):

Anytime =)